### v2

# CRAFT: Continuous Reasoning and Agentic Feedback Tuning for Multimodal Text-to-Image Generation

Abstract
Recent work has shown that inference-time reasoning and reflection can improve text-to-image generation without retraining. However, existing approaches often rely on implicit, holistic critiques or unconstrained prompt rewrites, making their behavior difficult to interpret, control, or stop reliably. In contrast, large language models have benefited from explicit, structured forms of thinking based on verification, targeted correction, and early stopping.
We introduce CRAFT (Continuous Reasoning and Agentic Feedback Tuning), a training-free, model-agnostic framework that brings this structured reasoning paradigm to multimodal image generation. CRAFT decomposes a prompt into dependency-structured visual questions, verifies generated images using a vision-language model, and applies targeted prompt edits through an LLM agent only where constraints fail. The process iterates with an explicit stopping criterion once all constraints are satisfied, yielding an interpretable and controllable inference-time refinement loop.
Across multiple model families and challenging benchmarks, CRAFT consistently improves compositional accuracy, text rendering, and preference-based evaluations, with particularly strong gains for lightweight generators. Importantly, these improvements incur only a negligible inference-time overhead, allowing smaller or cheaper models to approach the quality of substantially more expensive systems. Our results suggest that explicitly structured, constraint-driven inference-time reasoning is a key ingredient for improving the reliability of multimodal generative models.
1 Introduction
Recent progress in large language models has demonstrated that explicit inference-time reasoning–including verification, reflection, and targeted correction–can substantially improve reliability without additional training. Rather than relying solely on larger models or additional data, these systems benefit from structured forms of thinking that make intermediate decisions explicit, checkable, and correctable. Related ideas have recently begun to influence multimodal generation, where reflection-based and agentic pipelines show that post-hoc critique can improve text-to-image (T2I) outputs without retraining [19, 16].
At the same time, modern T2I systems remain predominantly single-shot and reactive. Despite impressive visual fidelity achieved by diffusion- and transformer-based models [14, 13, 15, 11, 1, 18], they often struggle with long, compositional, and text-heavy prompts, where multiple constraints must be satisfied simultaneously. Generation is typically controlled through a single textual channel–the prompt–which acts as the primary mechanism for steering content, composition, style, and even safety in black-box deployments.
While inference-time refinement can mitigate some of these issues, existing approaches often rely on implicit, holistic critiques or unconstrained prompt rewrites [12, 19, 16], making their behavior difficult to interpret, control, or stop reliably. A key challenge is that image quality in open-domain T2I is inherently multi-criteria: a satisfactory image must satisfy compositional constraints, maintain object attributes and relations, adhere to style and layout, render text correctly, and avoid visual artifacts–all at once.
Recent work therefore emphasizes explicit, question-based evaluation, where prompts are decomposed into verifiable visual statements that better correlate with human judgments [7, 8]. Benchmarks such as Davidsonian Scene Graph (DSG) and GenEval further demonstrate that structured, checkable constraints provide a more reliable foundation for evaluating and improving T2I systems [4, 5]. These insights suggest that open-domain T2I models could similarly benefit from converting implicit textual requirements into explicit, verifiable questions.
2 Related Work
We propose CRAFT, a training-free reasoning layer for text-to-image generation and image editing (Figure 1) that performs iterative refinement while avoiding redundant steps once the image satisfies the target quality criteria. In our implementation, the text-to-image and image-editing components of CRAFT can be used either sequentially or independently, which allows the framework to be integrated flexibly into diverse generation pipelines and model stacks.
Figure 1: Schema of CRAFT image generation and image editing.
3 CRAFT T2I Generator
The generator works iteratively. Its goal is to turn the initial prompt into a controlled, unambiguous improved prompt and a best image under the DVQ criterion.
Figure 2: Schema of CRAFT T2I generator.
Workflow.
Initialization. From the user prompt $\mathcal{P}_{0}$, two processes start in parallel:
Visual-question generation: We construct a set of explicit visual questions $\mathcal{Q}$ that describe the intended scene and, when applicable, a dependency graph $\mathcal{G}$ so that prerequisite questions precede their dependents (e.g., “Is there a cat?” before “Is the cat red?”) [4]. The base question set covers objects, attributes, spatial relations, and high-level scene properties derived from the user prompt. In addition, we augment $\mathcal{Q}$ with two generic diagnostic queries: a text-visibility question “Is the text in the attached image fully clear and easily readable by a human?” and an artifact check “Does the image look natural and free from visual glitches, artifacts, or distortions?” Both are answered strictly with YES/NO to keep the signal discrete. The text-visibility question is included only when the prompt contains explicit textual content that is intended to appear in the image (e.g., spans in quotation marks or other text-to-render instructions). In such cases, the question acts as a global check on whether the rendered typography is legible, complementing the fine-grained DVQ questions about whether specific strings appear. The artifact check is always present and serves as a generic quality probe, capturing issues such as distorted text, warped geometry, or rendering glitches. Together, these additions extend the question set beyond purely semantic correctness to cover text legibility and low-level visual fidelity, which are then used as constraints in subsequent refinement steps.
Prompt rewrite: use an LLM to rewrite $\mathcal{P}_{0}$ following best practices for T2I prompting, reducing ambiguity while preserving the original intent [12]. At each iteration $t$, the system applies a controlled rewriting step to produce an improved prompt while preserving the user’s intent. The rewriting module is implemented as a constrained LLM-based editor that operates under a fixed set of rules. First, the module enforces a canonical structure for text-to-image prompting – explicitly ensuring that the rewritten prompt contains a well-formed subject description, contextual/environmental details, and a stylistic specification. If the original prompt does not specify a style, the model inserts a default realistic style. The editor also enhances semantic clarity by adding descriptive modifiers and resolving ambiguous phrasing, while avoiding stylistic drift beyond the bounds of the user’s initial intent.
A key component of the rewriting logic is the explicit handling of textual elements. If the original prompt contains text enclosed in quotation marks (e.g., headers, labels, signage), this text is treated as literal content that must appear verbatim in the generated image. The rewriting module preserves such spans exactly – including case, punctuation, and quotation marks – and blocks their rephrasing, translation, or removal. In contrast, unquoted textual references (e.g., “with the title XX”, “named YY”) are interpreted as metadata or thematic context rather than visible text; these are allowed to be paraphrased or reframed as stylistic descriptors. This distinction prevents both unintended text hallucination and accidental loss or corruption of required text elements.
Generation. Using the updated prompt $\mathcal{P}_{t}$, the system generates an image $\mathcal{I}_{t}$ with a text-to-image (T2I) model $\mathcal{M}_{\text{T2I}}$:
$\mathcal{I}_{t}=\mathcal{M}_{\text{T2I}}(\mathcal{P}_{t};\theta_{t})$ (1)
where $\theta_{t}$ represents deterministic sampling parameters tied to iteration index. The framework is model-agnostic: any black-box T2I generator can be plugged into this stage without architectural changes, including diffusion-based models (e.g., FLUX, SD derivatives), autoregressive generators, or proprietary APIs.
[…]
$\text{score}_{\text{DVQ}}(\mathcal{I}_{t})=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathbb{1}[\mathcal{M}_{\text{VLM}}(\mathcal{I}_{t},q)=\text{YES}]$ (2)
These rationales guide the next prompt edits [19] for better generation results.
Selection (iterative update).
  • First iteration: skip comparison and set the incumbent best to the current image and prompt; store the DVQ answers and rationales.
  • Later iterations: compare the newly generated image against the incumbent using the DVQ score $\text{score}_{\text{DVQ}}(\mathcal{I}_{t})$. If the new image yields a higher score, it becomes the new incumbent:
    $\mathcal{I}_{\text{best}}=\begin{cases}\mathcal{I}_{t}&\text{if }\text{score}_{\text{DVQ}}(\mathcal{I}_{t})>\text{score}_{\text{DVQ}}(\mathcal{I}_{\text{best}})\\ \text{Judge}(\mathcal{I}_{t},\mathcal{I}_{\text{best}},\mathcal{P}_{0})&\text{if scores are equal}\end{cases}$ (3)
    When the YES counts are equal, we invoke an MLLM-based side-by-side comparison. The judge selects the image that better satisfies the prompt and exhibits fewer visual artifacts.
Targeted prompt update. Provide the initial prompt $\mathcal{P}_{0}$, the current prompt $\mathcal{P}_{t}$, the question set $\mathcal{Q}$, and all DVQ answers and rationales to an LLM $\mathcal{M}_{\text{LLM}}$. The LLM produces a targeted revision:
$\mathcal{P}_{t+1}=\mathcal{M}_{\text{LLM}}(\mathcal{P}_{0},\mathcal{P}_{t},\mathcal{Q},\{r_{q}:q\in\mathcal{Q}_{\text{failed}}\})$ (4)
that aims to increase the YES rate on $\mathcal{Q}$ while preserving attributes from $\mathcal{P}_{0}$.
Stopping. Run for a configurable maximum number of iterations $T_{\max}$, or stop early when:
$\text{score}_{\text{DVQ}}(\mathcal{I}_{t})=1.0\quad\text{(all questions answered YES)}$ (5)
4 CRAFT Image Editing
CRAFT iterative image editing (Figure 3) follows the same procedure as the iterative image generation loop, with two adaptations: (i) the prompt-rewriting and task prompts are tailored for editing, and (ii) the inputs consist of both a prompt and an image. The module can run independently or be composed with the iterative image generator.
Figure 3: CRAFT image editing schema.
5 Dataset and Models
We evaluated using DSG and DVQ metrics [4]. In addition to the main 2-iteration configuration, we tested multiple backbone models and multiple iteration budgets for both generation and editing (see Model configurations for baseline and CRAFT). For each configuration, we computed all metrics across several datasets to assess robustness and cross-model consistency. For text-to-image (T2I), we ran a 2-iteration setup with the Qwen-Image model [17]. For image editing, we used a 2-iteration setup with Gemini 2.5 Flash by Google. We further included two preference-based metrics inspired by the Maestro paper – auto side-by-side and auto side-by-side advantage [16]. As evaluation data, we used DSG-1K [4] and the Parti-Prompt (P2) suite [18]. We also report results on the P2-hard subset of Parti-Prompt, following the Maestro protocol. Results are presented below (Table 1, Table 3).
5.1 Model Configurations for Baseline and CRAFT
For our evaluations, we consider five families of T2I models. For each model, we report results in two configurations: a baseline setting and a corresponding CRAFT setting.
Baseline 1 (B1) – FLUX-Schnell:
Images are generated with FLUX-Schnell using a single forward pass and standard generation parameters.
CRAFT 1 (C1) – FLUX-Schnell (2 iterations, no image editing):
We apply the full CRAFT pipeline for two iterations, using FLUX-Schnell for all generations, without any image-editing model. Only prompt-level improvements and DVQ-guided refinement are used.
Baseline 2 (B2) – FLUX-Dev:
Images are generated with FLUX-Dev using a single forward pass and standard generation parameters.
CRAFT 2 (C2) – FLUX-Dev (2 iterations, no image editing):
We apply the full CRAFT pipeline for two iterations, using FLUX-Dev for all generations, without any image-editing model.
Baseline 3 (B3) – Qwen-Image (1 iteration):
Images are generated with Qwen-Image using default settings, without iterative refinement.
CRAFT 3 (C3) – Qwen-Image (2 T2I iterations) + Gemini 2.5 Flash (2 editing iterations):
We run two CRAFT iterations with Qwen-Image, followed by two DVQ-guided image-editing iterations using Gemini 2.5 Flash.
Baseline 4 (B4) – Z-Image-Turbo (1 iteration):
Images are generated with Z-Image-Turbo using default settings and a single forward pass.
CRAFT 4 (C4) – Z-Image-Turbo (2 T2I iterations):
We run two DVQ-guided CRAFT iterations with Z-Image-Turbo.
Baseline 5 (B5) – FLUX-2 Pro (1 iteration):
Images are generated with FLUX-2 Pro using default parameters.
CRAFT 5 (C5) – FLUX-2 Pro (2 T2I iterations):
We run two CRAFT iterations with FLUX-2 Pro, using DVQ feedback to iteratively refine the prompt.
6 Implementation Details
This section describes the specific models, hyperparameters, and system configurations used in our CRAFT implementation.
6.1 Model Selection
Vision-Language Model (VLM):
All VQA evaluation, DVQ question answering, image critique, and side-by-side comparison steps use GPT-5-nano (gpt-5-nano-2025-08-07) via OpenAI API. We selected this model for its balance of cost efficiency, fast inference speed, and strong vision-understanding capabilities. The model demonstrates robust performance on binary visual question answering tasks and provides consistent, structured outputs suitable for automated evaluation pipelines.
Language Model (LLM):
Prompt rewriting, DVQ question generation, and targeted refinement operations use GPT-5-nano (gpt-5-nano-2025-08-07). The model’s strong instruction-following capabilities and ability to maintain semantic consistency across iterative edits make it well-suited for constrained prompt engineering tasks.
Text-to-Image Backbones:
We evaluate CRAFT with five T2I models representing different performance and cost tiers:
  • FLUX-Schnell and FLUX-Dev: Open-source diffusion models with strong baseline quality [2].
  • Qwen-Image: Multimodal autoregressive model with native text-to-image generation [17].
  • Z-Image-Turbo: Novelly published, open-source diffusion model with fast inference speed and one of the best baseline quality  [3].
  • FLUX-2 Pro: Commercial API with enhanced text rendering and compositional capabilities. All models are accessed via their respective APIs with default parameters unless otherwise specified in ablation studies.
Image Editing Model:
For CRAFT image editing experiments, we use Gemini 2.5 Flash by Google for its strong instruction-following in image manipulation tasks and efficient inference.
6.2 Hyperparameters and Configuration
Temperature and Sampling:
All LLM and VLM calls use temperature=0.0 for deterministic, reproducible outputs. This ensures consistency across runs and eliminates sampling variance in evaluation metrics.
Generation Parameters:
For reproducibility, we use fixed seeds per prompt across all baseline and CRAFT configurations. T2I model parameters (guidance scale, inference steps) follow default API settings for each backbone to ensure fair comparison with typical user experiences. Questions used for DSG and DVQ metrics are generated using once for baseline prompt and use for evaluate in both baseline and CRAFT.
6.3 System Prompts
Each component of the CRAFT pipeline uses a carefully designed system prompt to guide the LLM or VLM behavior:
  • Initial Prompt Rewriting: Transforms user prompts into T2I-optimized format while preserving intent (Appendix A.1).
  • Verify and Self Correct: Verify whether the original prompt satisfies the constraints specified by the user and self-correct the prompt if necessary (Appendix A.2).
  • DSG Question Generation: Constructs visual questions and dependency graphs from prompts, as it was implemented in [4].
  • VQA Evaluation: Performs binary YES/NO visual question answering with rationales (Appendix A.3).
  • DVQ Rationalization: Rationalize the DVQ answers with answer "NO" to the image (Appendix A.4).
  • Targeted Prompt Refinement: Edits prompts based on DVQ feedback to address specific failures (Appendix A.5).
  • Side-by-Side Judge: Compares two images to select the better match for a given prompt (Appendix A.6).
  • Image Editing: Applies targeted modifications to images based on structured instructions (Appendix A.7).
Full prompt templates are provided in Appendix A to ensure reproducibility.
6.4 Computational Cost
The additional inference-time cost introduced by CRAFT reasoning is negligible compared to T2I generation costs. For a typical $2048\times 2048$ image with 2 iterations:
  • VLM cost per iteration: $\approx$\$0.0012 (vision tokens + text output for $2048\times 2048$ images)
  • LLM cost per iteration: $\approx$\$0.00001 (prompt rewriting)
  • Total CRAFT overhead: $\approx$\$0.0024 for 2 iterations
This represents approximately 0.6% of the cost of generating a single $2048\times 2048$ image with premium T2I models like Hunyuan Image 3.0 (\$0.42 per image). The cost analysis is detailed further in Section 9.
Latency.
A single CRAFT iteration (generation or editing) takes $\approx 30$ seconds on average in our setup; three iterations typically yield strong results. In contrast, a single Hunyuan Image 3.0 generation averages $\approx 60$ seconds while underperforming CRAFT-equipped mid-cost models on compositional quality.
7 Metrics
7.1 VQA (DVQ / Visual Question Answering)
The VQA-based metric evaluates an image by asking a multimodal LLM a set of deterministic, unambiguous visual questions derived from the prompt. The model answers each question with a strict “yes” or “no.” The resulting yes-rate serves as a direct diagnostic score indicating how well the image matches each explicit constraint in the prompt:
$\text{score}_{\text{VQA}}(\mathcal{I})=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathbb{1}[\text{answer}_{q}=\text{YES}]$ (6)
where $\mathcal{Q}$ is the set of visual questions and $\mathcal{I}$ is the generated image.
7.2 DSG (Davidsonian Scene Graph)
DSG (Davidsonian Scene Graph) measures prompt–image consistency by converting a prompt into a set of structured, binary visual questions. These questions are arranged as a dependency graph $\mathcal{G}$, where a question is evaluated only if its parent conditions hold (for example, an attribute such as color is checked only after the corresponding object is confirmed to exist). The DSG score is computed as:
$\text{score}_{\text{DSG}}(\mathcal{I})=\frac{\sum_{q\in\mathcal{Q}_{\text{valid}}}\mathbb{1}[\text{answer}_{q}=\text{YES}]}{|\mathcal{Q}_{\text{valid}}|}$ (7)
where $\mathcal{Q}_{\text{valid}}$ is the set of questions whose parent conditions are satisfied, yielding a compositional, hierarchy-aware estimate of image faithfulness.
7.3 Auto Side-by-Side
Auto Side-by-Side is an automatic preference metric inspired by human A/B evaluations. A multimodal LLM is shown two images $\mathcal{I}_{\text{baseline}}$ and $\mathcal{I}_{\text{CRAFT}}$ along with the original prompt $\mathcal{P}$ and asked to select which image better satisfies the instruction. The win-rate is computed as:
$\text{win-rate}_{\text{CRAFT}}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}[\text{Judge}(\mathcal{I}_{i}^{\text{CRAFT}},\mathcal{I}_{i}^{\text{baseline}},\mathcal{P}_{i})=\mathcal{I}_{i}^{\text{CRAFT}}]$ (8)
where $N$ is the total number of test prompts. This produces a binary preference score that correlates with human judgments while remaining model-agnostic.
8 Results
8.1 Results on DSG1K Dataset
Table 1 demonstrates that CRAFT yields consistent and substantial improvements across all evaluated model backbones on the DSG1K benchmark. For every configuration – ranging from lighter diffusion models such as FLUX-Schnell and FLUX-Dev to more advanced systems like Qwen-Image, Z-Image-Turbo, and FLUX-2 Pro – the iterative refinement pipeline increases both VQA and DSG scores and leads to pronounced gains in preference-based metrics.
Table 1: Results on DSG1K [4] dataset.
Metric B1 C1 B2 C2 B3 C3 B4 C4 B5 C5
VQA 0.78 0.86 0.77 0.87 0.864 0.946 0.875 0.915 0.88 0.925
DSG 0.786 0.857 0.765 0.841 0.858 0.932 0.864 0.9 0.87 0.91
Auto SxS 0.21 0.744 0.34 0.65 0.36 0.59 0.27 0.67 0.4 0.52
Auto SxS Adv 0.26 0.74 0.288 0.671 0.33 0.58 0.328 0.63 0.47 0.5
8.2 Comparison with Prompt Optimization Methods
Table 2 compares CRAFT with prompt-optimization baselines reported in Maestro [16]. We report only DSGScore and omit ranks. Note that our CRAFT results use a GPT-based VLM judge (gpt-5-nano-2025-08-07), whereas Maestro reports scores with Gemini 2.0 Flash, so absolute values are not directly comparable across papers.
Table 2: Comparison with prompt-optimization methods from Maestro. We report DSGScore only; higher is better.
Method P2-hard DSGScore $\uparrow$ DSG-1K DSGScore $\uparrow$
Original 0.826 0.772
Rewrite 0.855 0.815
Promptist [6] 0.873 0.849
LM-BBO [9] 0.859 0.806
OPT2I [10] 0.900 0.838
Maestro [16] 0.92 0.882
CRAFT (Ours) 0.90 0.91
CRAFT achieves the best DSGScore on DSG-1K and is competitive on P2-hard, where Maestro reports slightly higher scores. Overall, constraint-driven refinement provides stronger compositional gains than generic prompt-rewrite or agentic optimization methods.
Notably, models with lower baseline compositional reliability benefit the most: for example, FLUX-Dev and FLUX-Schnell exhibit improvements of 881010 percentage points in VQA and DSG, with Auto SxS win-rates more than doubling (Figure 4). At the same time, even stronger models such as Qwen-Image, Z-Image-Turbo, and FLUX-2 Pro continue to improve, indicating that the proposed approach scales with model capacity and is not limited to weaker backbones.
Figure 4: Auto SxS win rates on DSG1K. CRAFT is preferred over baseline across all model configurations.
8.3 Results on Parti-Prompt Dataset
Table 3 shows that the proposed CRAFT leads to consistent improvements across all evaluated model families on the Parti-Prompts dataset. For lighter T2I backbones such as FLUX-Dev and FLUX-Schnell, two reasoning iterations yield clear gains in both VQA and DSG scores, often closing much of the gap to stronger baselines. Higher-capacity models such as Qwen-Image and FLUX-2-Pro also benefit from iterative refinement, though their improvements are smaller in absolute magnitude due to already strong baseline performance.
Table 3: Results on Parti-Prompt dataset.
Metric B1 C1 B2 C2 B3 C3 B4 C4 B5 C5
VQA 0.79 0.88 0.822 0.911 0.89 0.94 0.9 0.94 0.90 0.92
DSG 0.743 0.857 0.815 0.895 0.896 0.939 0.89 0.92 0.87 0.90
Auto SxS 0.19 0.756 0.382 0.618 0.445 0.555 0.268 0.679 0.42 0.47
Auto SxS Adv 0.252 0.748 0.309 0.628 0.41 0.53 0.328 0.668 0.47 0.52
Preference-based metrics (Auto SxS and Auto SxS advantage) indicate that CRAFT is systematically preferred over baseline generations for most models (Figure 5), with particularly large shifts for weaker backbones. Qualitative comparisons further align with these quantitative trends: across all architectures, CRAFT reduces omissions, improves attribute consistency, and produces more coherent text rendering and layout for text-heavy prompts.
Figure 5: Auto SxS win rates on Parti-Prompt. CRAFT is preferred over baseline across all model configurations.
8.4 Ablation: Effect of Number of Iterations
Figure 6: VQA and DSG scores as a function of the number of iterations for CRAFT with FLUX-Schnell.
Figure 6 demonstrates that the introduction of even a single reasoning iteration yields a substantial improvement in both VQA and DSG scores relative to the one-shot FLUX baseline. Additional iterations continue to provide incremental gains, though with diminishing returns, and importantly do not degrade performance. This behavior suggests that the “thinking” loop is most effective in its first pass – correcting major prompt–image inconsistencies – while subsequent iterations refine remaining details without destabilizing generation quality.
8.5 Text Rendering Quality: Normalized Edit Distance
We evaluate text rendering quality using Normalized Edit Distance (NED), computed as:
$\text{NED}(x,y)=\frac{\text{Levenshtein}(x,y)}{\max(|x|,|y|)}$ (9)
where $\text{Levenshtein}(x,y)$ measures character-level edits (insertions, deletions, substitutions) and normalization ensures fair comparison across different phrase lengths. For each prompt, we extract expected text phrases using an LLM and perform OCR on generated images using a Vision-Language Model. We report Text Accuracy $1-\text{NED}$ for intuitive interpretation and Exact Match Rate (percentage of phrases with $\text{NED}=0$) as a strict quality measure.
Table 4 shows that CRAFT yields measurable improvements in text rendering performance for models with weaker baseline text fidelity. FLUX-Schnell exhibits the largest gains in both Text Accuracy and Exact Match rate, indicating that iterative constraint-based refinement reduces character-level deviations. FLUX-Dev shows moderate improvements, consistent with its baseline limitations. For Z-Image and FLUX-2 Pro, changes introduced by CRAFT are minor, reflecting their relatively stable baseline behavior and limited headroom for improvement in this task.
Table 4: Text rendering quality evaluated using Normalized Edit Distance (NED).
Model Baseline Acc. CRAFT Acc. Improvement Exact Match Gain
FLUX-Schnell 74.01% 81.17% +7.16% +12.91%
FLUX-Dev 64.01% 68.90% +4.89% +6.00%
Z-Image 91.00% 91.40% +0.40% +1.10%
FLUX-2 Pro 98.50% 98.80% +0.30% +0.25%
8.6 Human Evaluation
To complement the automatic preference metrics, we conducted an additional small-scale human evaluation on two model families: Z-Image-Turbo and FLUX-2 Pro. Twenty evaluators (who work in the field of AI content generation) were shown two images (baseline and CRAFT mode) and asked to select the version that best satisfied the original prompt. Ties were allowed.
Across both models, the results closely mirrored the automatic side-by-side metrics, consistently favoring the CRAFT outputs (Figures 710). These results indicate that the improvements observed under DVQ and side-by-side metrics are also reflected in human preference judgments.
Figure 7: Human evaluation results for FLUX-Schnell model.
Figure 8: Human evaluation results for FLUX-Dev model.
Figure 9: Human evaluation results for Z-Image-Turbo model.
Figure 10: Human evaluation results for FLUX-2 Pro model.
8.7 Qualitative Results
We present qualitative comparisons across all evaluated models. Across backbones, reasoning-guided refinement yields improved attribute consistency, reduced omissions, more coherent spatial relations, and better handling of long prompts with multiple clauses or textual regions.
8.7.1 Qwen Image 2 iterations + Gemini Flash 2.5
Figure 11: Baseline vs CRAFT. Prompt: A giant cobra snake made from salad.
Figure 12: Baseline vs CRAFT. Prompt: a toaster shaking hands with a microwave.
Figure 13: Baseline vs CRAFT. Prompt: a view of the Earth from the moon.
We also demonstrate results for long-form prompts and for prompts that include text elements rendered within the generated images (see Appendix B for full prompt listings).
Figure 14: CRAFT results for long_prompt1.
8.7.2 FLUX-Dev 2 iterations
We also present a comparison between the baseline and the CRAFT (2 iterations) for FLUX-Dev, without using any image-editing models. All generations were produced with standard parameters. Even this lightweight setup – simply correcting the prompt based on the VQ questions and applying the thinking procedure – already leads to noticeable improvements in generation quality. Additional examples are provided in Appendix D.
Figure 15: Baseline vs CRAFT. Prompt: A woman holding a racquet on top of a tennis court.
Figure 16: Baseline vs CRAFT. Prompt: studio shot multicolored fur in the shape of word ’hello’, in a furry frame, white background, centered.
Figure 17: Baseline vs CRAFT. Prompt: A woman’s mirror reflection is wearing different clothes.
8.7.3 Z-Image 2 iterations
We additionally evaluate Z-Image under a two-iteration CRAFT, without any image-editing components. Additional examples are provided in Appendix D.
Figure 18: Baseline vs CRAFT. Prompt: Generate a series of six candid, documentary-style photos of this Indonesian president in office, in the rice fields, and partying with other presidents.
Figure 19: Baseline vs CRAFT. Prompt: Generate a 4‑panel comic about the hardships of an embedded engineer.
8.7.4 FLUX-2 Pro 2 iterations
We additionally evaluate FLUX2-Pro under a two-iteration CRAFT, without any image-editing components. Additional examples are provided in Appendix D.
Figure 20: Baseline vs CRAFT. Prompt: Dream diary. A pink Kirby lying asleep on a star, blowing rainbow-colored bubbles from its mouth. Soft macaron color palette, cloud and candy stickers, glittery crayon-like details, dreamy and sweet.
Figure 21: Baseline vs CRAFT. Prompt: Create a technical exploded view diagram of an iPhone pro max in 1080x1080 dimensions. The device should be deconstructed, showing all individual components floating in space, separated to reveal the internal parts. Each major component is labeled.
We observed that some prompts in the PartiPrompt dataset produce unreliable generations: several entries consist of only one or two highly abstract words, which lead to poor results in both the baseline and the CRAFT. Examples include prompts such as “emotion” or “hate”. To address this, we filtered the PartiPrompt subset and retained only prompts containing more than three words.
Figure 22: CRAFT vs baseline for prompt: intelligence.
Figure 23: Baseline vs CRAFT for prompt: commonsense.
We conducted evaluations using GenEval, but the results were not representative, as the available detectors performed poorly even on clear, high-quality cases. The detectors used were those provided in MMCV 1.7.1, including QueryInst R101, Mask2Former, and YOLOX-L.
Figure 24: CRAFT versus baseline for prompt: a photo of a keyboard.
9 Cost and Quality Comparison
To complement the compositional and preference-based evaluations, we analyze the cost–quality trade-off of applying CRAFT to lightweight T2I generators. Table 5 presents a detailed breakdown of generation costs and quality gains at a fixed resolution of $2048\times 2048$ pixels ($\approx 4.19$ megapixels) across five backbone models.
9.1 Cost Model
For a backbone with price $c$ dollars per megapixel (MP), the cost of generating a single image at resolution $r$ MP is $c\cdot r$. CRAFT with $T$ iterations requires $T$ generation calls plus multimodal reasoning overhead. For T=2T{=}2 iterations at r=4.19r{=}4.19 MP, the total cost becomes:
$\text{Cost}_{\text{CRAFT}}=2\cdot c\cdot r+\text{Cost}_{\text{reasoning}}$ (10)
where $\text{Cost}_{\text{reasoning}}\approx\mathdollar 0.0024$ (two VLM evaluation cycles using GPT-5-nano for $2048\times 2048$ images).
9.2 Empirical Cost–Quality Analysis
Table 5 reveals four key patterns:
Mid-cost generators achieve premium quality below premium cost.
Z-Image-Turbo (\$0.005/MP) with CRAFT costs $\approx\mathdollar 0.044$ per image—over $9.5\times$ cheaper than a single Hunyuan Image 3.0 call (\$0.42)—while substantially improving quality (DSG1K: VQA $4.6\%$, DSG $4.2\%$, Auto SxS win rate +40.0+40.0 $p.$).
Low-cost generators see largest relative gains.
FLUX-Schnell (\$0.003/MP) achieves the largest quality improvements (VQA $10.3\%$, DSG $9.0\%$, Auto SxS +53.4+53.4 p.p.) at minimal absolute cost ($\approx\mathdollar 0.029$ total, $2.2\times$ baseline). Here, reasoning overhead ($\approx 8.3\%$ of total cost) becomes more visible, suggesting adaptive early stopping or cheaper verifiers would be particularly beneficial.
Reasoning overhead is negligible for mid-to-high-cost generators.
For FLUX-Dev (\$0.025/MP), Z-Image-Turbo (\$0.005/MP), and FLUX-2 Pro (\$0.0179/MP), the fixed reasoning cost represents only $1.1\%$, $5.5\%$, and $1.6\%$ of total CRAFT cost, respectively, making the trade-off highly favorable.
Premium models benefit from targeted refinement.
FLUX-2 Pro (\$0.0179/MP), despite having strong baseline performance, shows consistent improvements through CRAFT (VQA $5.1\%$, DSG $4.6\%$, Auto SxS +12.0+12.0 p.p.). At $\approx\mathdollar 0.152$ total cost ($2.0\times$ baseline), it delivers state-of-the-art quality while remaining $\approx 64\%$ cheaper than a single Hunyuan Image 3.0 call.
Latency and iteration budget.
In our experiments, a single CRAFT iteration (generation or editing) takes on average $\approx 30$ seconds, and three iterations are typically sufficient to reach strong qualitative results. By comparison, a single Hunyuan Image 3.0 generation takes $\approx 60$ seconds on average while delivering lower compositional quality than CRAFT-equipped mid-cost backbones under our evaluation.
Table 5: Cost–quality trade-off at $2048{\times}2048$ resolution. CRAFT cost includes two generations plus \$0.0024 reasoning overhead. Cost multiplier shown in parentheses. Quality gains on DSG1K: VQA and DSG as relative improvement (%); Auto SxS as absolute difference in percentage points ($p.$). Best gains per column in bold.
Backbone $\mathdollar/\text{MP}$ Cost @ $2048{\times}2048$ Quality Gain
[…]
Reference (no CRAFT applied):
Hunyuan 3.0 0.1000 \$0.42
9.3 Qualitative Comparison
Representative visual comparisons (Figures 2526) between Z-Image-Turbo baseline, Z-Image-Turbo with CRAFT, and Hunyuan Image 3.0 demonstrate that reasoning-guided refinement substantially improves object–attribute correctness and spatial relationships, often narrowing the gap between lightweight and heavyweight models. Full prompts are provided in Appendix C. Additional examples are available in Appendix E.
Figure 25: Baseline vs CRAFT vs Hunyuan Image 3 (cost_prompt1, Appendix C).
Figure 26: Baseline vs CRAFT vs Hunyuan Image 3 (cost_prompt2, Appendix C).
10 Limitations
Despite its generality, the proposed CRAFT has several limitations. First, the refinement loop depends on the reliability of the multimodal LLM responsible for visual question evaluation, text recognition, and side-by-side judgment. Although stronger MLLMs reduce noise, failures in text reading, misinterpretation of fine-grained attributes, or aesthetic biases can propagate incorrect supervisory signals across iterations. This is particularly evident for prompts with dense textual content, where OCR errors or hallucinated legibility assessments may mislead the update step.
Second, while the DVQ framework covers compositional, textual, and artifact-related constraints, it still requires that the prompt be mappable to a well-defined question set $\mathcal{Q}$. Highly abstract or underspecified prompts – commonly present in datasets such as Parti-Prompt – remain challenging, even after filtering, because they lack concrete visual targets for iterative correction.
Third, inference-time cost remains higher than single-shot generation: each iteration requires multiple MLLM and T2I calls. Although most gains materialize within one or two iterations, the overhead may be non-negligible for large-scale or low-latency deployments.
Fourth, evaluation of text rendering quality – while improved via NED – remains limited by OCR fidelity on stylized or specific-language non-Latin typography.
Finally, while human evaluations (Section 8.6) confirm the preference trends observed with automatic side-by-side metrics, the study remains small-scale.
11 Conclusion
We introduced CRAFT, a training-free and model-agnostic reasoning framework that improves text-to-image and image-editing systems through explicit, constraint-driven verification and targeted prompt refinement at inference time. Across all evaluated architectures–from lightweight generators to high-end diffusion models–the method consistently enhances compositional accuracy, textual fidelity, artifact detection, and multimodal preference behavior. Notably, weaker backbones often approach the performance of substantially stronger models after only one or two reasoning iterations, indicating that structured inference-time reasoning can partially compensate for architectural limitations without retraining.
A central advantage of the framework is its full modularity. CRAFT operates as a plug-in layer that can be paired with any text-to-image generator, any vision–language model for verification, and any large language model for prompt rewriting, without model-specific adaptation. As a result, the system directly benefits from improvements in underlying components, supports heterogeneous generation pipelines, and can be integrated with minimal engineering effort in black-box deployment settings.
Because the approach relies on explicit verification, targeted correction, and a clear stopping criterion, it is not inherently restricted to the image domain. The same principles suggest promising directions for extending structured, constraint-driven inference-time reasoning to other multimodal generative settings, such as video generation and editing (via temporal constraints) or interactive content generation, provided that intermediate outputs can be evaluated and corrected in a principled manner.
From a practical standpoint, the reasoning loop introduces negligible cost overhead relative to modern high-resolution text-to-image models. For example, running two reasoning iterations for Z-Image-Turbo costs approximately \$$0.0024$ per image, corresponding to less than $0.3\%$ of the generation cost of Hunyuan Image 3.0 at comparable resolution. Despite operating at more than $2\times$ lower cost per megapixel, Z-Image-Turbo equipped with CRAFT often matches or exceeds the compositional performance of Hunyuan Image 3.0 in our evaluations, demonstrating that inference-time reasoning offers a cost-efficient alternative to scaling model size or relying on more expensive diffusion backbones.
Overall, our results show that inference-time reasoning becomes reliable and scalable when it is made explicit, structured, and verifiable. By grounding refinement in constraint satisfaction rather than opaque holistic scores, CRAFT provides a practical and interpretable path toward improving prompt adherence and reducing failure modes in multimodal generative systems. We hope this work encourages further exploration of structured reasoning paradigms as a foundation for more controllable and dependable generative models.
Appendix A System Prompts
This section provides the complete system prompts used for each component of the CRAFT pipeline. All prompts use GPT-5-nano (gpt-5-nano-2025-08-07) with temperature=0.0 for reproducibility.
A.1 Initial Prompt Rewriting
System Prompt:
I want to create an image by giving a prompt to a text-to-image model.
Here is my current prompt: $\{$ current\_prompt \}
Please generate an improved prompt based on the following guidelines for text-to-image generator:
1. Make sure the prompt contains the elements of subject, context, and style. Specifically:
- Subject: The main object, person, animal, or scenery you want an image of.
- Context and background: Also consider the background or environment for the subject. For example, a studio, outdoor scenery, or an indoor setting.
- Style: Add the style of image you want. Styles can be general (painting, photograph, sketches) or very specific (pastel painting, charcoal drawing, isometric 3D).
* IMPORTANT: If the original prompt does not explicitly specify a style, add "realistic style" by default in the improved prompt.
2. Use descriptive language: Employ detailed adjectives and adverbs to paint a clear picture for image generator.
3. Provide context: If necessary, include background information to aid the AIs understanding.
4. Reference specific artists or styles: If you have a particular aesthetic in mind, referencing specific artists or art movements can be helpful.
5. Enhancing facial details in images of people:
- Specify facial details as a focus of the photo (for example, use the word "portrait" in the prompt).
6. CRITICAL - Preserving text in quotation marks:
- If the original prompt contains text within quotation marks (e.g., "Hello World"), this is LITERAL TEXT that must appear in the generated image.
- You MUST preserve this text EXACTLY as written, including the quotation marks.
- DO NOT modify, rephrase, translate, or remove any text that appears in quotation marks.
- DO NOT remove the quotation marks themselves - they indicate visible text.
- Example: sign that says "Welcome"’ -> keep "Welcome" exactly as is with quotes
- Example: banner with text "Sale 50%"’ -> keep "Sale 50%" exactly as is with quotes
7. IMPORTANT - Handling titles vs visible text:
- If the prompt contains "with the title X", "titled X", "named X", or "called X" WITHOUT quotation marks - this refers to the ARTWORKS NAME, NOT text that should appear in the image.
- Rewrite these as style/theme descriptions without implying visible text.
- Example: "with the title Karakuri Ninja Oboro" -> "depicting a scene from Karakuri Ninja Oboro" or "featuring the Karakuri Ninja Oboro theme"
- Only text explicitly in quotation marks should be treated as visible text in the image.
8. Avoid ambiguity about text:
- Text WITH quotation marks = visible text in the image (preserve exactly)
- Text WITHOUT quotation marks = description/style/theme (can be rephrased)
- Never add quotation marks unless they were in the original prompt
- Never remove quotation marks if they were in the original prompt
Please now generate 1 improved prompt and enclose it between <PROMPT> and </PROMPT>.
If you believe the current prompt is already satisfactory, simply respond with "<PROMPT> NO\_CHANGE </PROMPT>".
A.2 Verify and Self Correct
System Prompt:
Your task is to first verify whether my original prompt satisfies each of the constraints specified by the user: $\{$ constraints \}
Make sure you verify each constraint and answer "YES" if the constraint is met or "NO" otherwise, followed by a short explanation.
If there is any unmet constraint, please revise the prompt to satisfy all constraints.
Make sure you only modify the part of the prompt that previously led to constraint violations and bracket any revised
prompt between <answer> and </answer>. If the original prompt satisfies all constraints, you may simply write
"<answer> NO_CHANGE </answer>".
A.3 VQA Evaluation
System Prompt:
Answer only with "Yes" or "No". Do not give other outputs or punctuation marks.
Image: $\{$ image \}
Question: $\{$ dvq[i] \}
A.4 DVQ Rationalization
System Prompt:
You are an expert in examining images generated by text-to-image generation models. The user specifies a list of questions, each
of which contains a property or characteristic that the generated image should have and the goal is to make sure that when an
reviewer reviews the generated image, they will answer "Yes" to each of these questions. However, the reviewer answered "No" to
the question below to the current image:
Image: $\{$ image \}
Question: $\{$ dvq[i] \}
A.5 Targeted Prompt Refinement
System Prompt:
You are an expert in prompt engineering for text-to-image generation models, which take a text prompt as input and generate images depicting the prompt as output. Your task is to improve upon my current prompt based on feedback I obtained from human reviewers.
The user has specified a list of questions, each of which contains a property or characteristic that the generated image should have and the goal is to make sure that when a reviewer reviews the generated image, they will answer "Yes" to all the questions. However, they have answered "No" to one or more questions and provided some feedback, and I’d like you to help me improve my existing prompt, taking into account their feedback.
My current prompt is: ’{best_prompt_so_far}’
But a reviewer looked at my image and answered "No" to the following question(s). Below are the questions and their feedback, which contains explanations why they believed the image does not satisfy the requirement in the question(s) and their suggestions for improvements:
{feedback_section}
Based on your analysis and the reviewer’s opinion, now improve my prompt by incorporating all their feedback points into new prompt(s) so that the reviewer will now answer "YES" to the question(s) when I generate a new image with the new prompt(s). Pay close attention to any points that they mentioned are lacking in the current image, and ensure you incorporate all these information in your revised prompts. Make sure the revised prompts are concise and effective. You may be creative in composing the new prompts, adding new elements, emphasize existing elements, or shuffle the order of the sentences to emphasize the previously missing characteristics at the beginning of the prompt. Finally, make sure each prompt you generated is bracketed between <PROMPT> and </PROMPT>
A.6 Side-by-Side Judge
System Prompt:
You are an expert in critiquing images. Given a user prompt, a text-to-image model generated two images
("Image A" and "Image B"), and your task is to pick the overall better image.
[…]
Please carefully explain the rationale before giving your final choice. Make sure you finish your answer with
"<answer> X </answer>", where X is "A" or "B".
A.7 Image Editing
System Prompt:
You are an expert in prompt engineering for image editing models, which take an instruction prompt and an existing image as input and produce an edited image as output. Your task is to improve my current editing instruction based on feedback I obtained from human reviewers.
The user has specified a list of questions, each describing a property or characteristic that the edited image should have. The goal is to ensure that, when a reviewer examines the edited image, they will answer ’Yes’ to all questions regarding these requirements. However, the reviewer answered "No" to one or more questions and provided feedback explaining why the image did not satisfy these requirements, along with suggestions for improvement.
My current image editing prompt is: ’{best_prompt_so_far}’
Below are the questions and the reviewer’s feedback, outlining what is lacking in the current edited image and any suggested improvements:
{feedback_section}
Based on your analysis and the reviewer’s feedback, improve my prompt by incorporating all their feedback and suggestions. Your revised editing instruction(s) should address every missing characteristic so that the reviewer will now answer ’Yes’ to every question when a new edit is made using your new prompt. Carefully consider all points mentioned as lacking and make sure they are explicitly incorporated in your improved prompts. Feel free to creatively rephrase, reorder, or emphasize specific elements to ensure clarity and effectiveness. Each improved instruction should be concise, clear, and actionable. Finally, enclose each improved prompt between <PROMPT> and </PROMPT>.
Appendix B Long-Form Prompts
B.1 long_prompt1
“Studio shot of modern cosmetic bottles on textured marble and velvet background, soft golden lighting, luxury minimalist style, 8K, hyperrealistic. Text elements: large header ‘LUMINOUS SKIN’ in a thin geometric sans-serif font (like Futura PT Light) in the top left corner, a central mathematical fraction ‘V/C = $\infty$’ in an elegant serif font (like Garamond) slightly off-center to the right, the caption ‘where Value meets Clarity’ in a thin italic serif font beneath the fraction, and a tagline ‘The future of radiance is here.’ in a light sans-serif font (like Montserrat Light) in the bottom right corner. Color palette: cream, charcoal, dark khaki, gold accents.”
B.2 long_prompt2
“Studio shot of a classic wooden chess set displayed in multiple angles on a clean white background, soft diffused daylight, product photography style, 8K, ultra-realistic. The main composition features a natural wood chessboard with alternating maple-light and walnut-dark squares, bordered by engraved coordinates in white. In the foreground, the board is fully set up for a game with traditional Staunton-style pieces in cream and ebony tones, arranged in perfect symmetry. Supporting shots above show the same board folded open (flat top view) and closed (vertical orientation), emphasizing craftsmanship and portability. Lighting is neutral and shadow-soft, highlighting the wood grain texture and subtle gloss finish. Text elements: Large minimalist header ‘CLASSIC CHESS SET’ in uppercase geometric sans-serif (e.g. Futura PT Medium) centered at the top; Sub-caption ‘Precision. Tradition. Balance.’ in light italic serif (e.g. Garamond Italic) beneath the header; Tagline ‘Handcrafted strategy for every mind.’ in thin sans-serif (e.g. Montserrat Light) in the lower right corner. Color palette: natural walnut brown, ivory, maple beige, white background, soft gray shadows. Mood: timeless elegance, craftsmanship, calm intellectual focus.”
B.3 long_prompt3
“Studio shot of a luxury single-watch travel case and timepiece on a seamless white background, minimalist product photography, 8K, ultra-realistic. The composition features a tan leather cylindrical case with soft suede interior, opened to reveal its plush texture and metal snap closures. Beside it, a stainless-steel wristwatch with a silver dial and black numerals rests on a matching suede-brown cushion, aligned precisely for balance and symmetry. Lighting: soft neutral daylight with gentle reflections on the watch bezel and bracelet, creating a clean, premium aesthetic. Shadows are soft and diffuse, emphasizing material contrast between metal, leather, and suede. Text elements: Large elegant header ‘TRAVEL IN TIME’ in all caps, geometric sans-serif (e.g. Futura PT Medium) at the top center; Sub-caption ‘A refined home for your watch.’ in light italic serif (e.g. Garamond Italic) below; Tagline ‘Crafted for journeys. Designed for care.’ in thin sans-serif (e.g. Montserrat Light) at the lower right corner. Color palette: warm caramel brown, beige suede, polished silver, white background, soft gray shadows. Mood: refined craftsmanship, modern classic elegance, understated luxury.”
B.4 long_prompt4
“Interior lifestyle shot of a modern bamboo laundry hamper in a minimalist home setting, 8K hyperrealistic, warm natural lighting. The hamper stands on a light wooden floor beside a cozy cream armchair draped with a knitted throw and a soft off-white rug. Its tall, rectangular form features smooth rounded edges, vertical bamboo slats, and a clean white fabric rim. A twisted rope handle adorns each side, matching the handle on the lid – blending natural materials with functional design. Lighting: soft daylight filtered through sheer curtains, creating gentle shadows and a serene, homely atmosphere. Composition: centered object focus, surrounded by neutral-toned decor – a wooden desk with a minimalist ceramic vase holding dried flowers, and a framed mirror subtly visible in the background. Color palette: warm honey bamboo, ivory white, beige, cream, and soft taupe accents. Mood: calm, clean, eco-conscious aesthetic – evoking Scandinavian simplicity and natural harmony. Text elements (optional for layout): Header: ‘Sustainable Living Starts at Home’ in soft sans-serif (e.g., Helvetica Neue Light); Caption: ‘Elegant bamboo design for everyday organization.’ in italic serif (e.g., Garamond Italic); Tagline: ‘Minimal form. Maximum calm.’ in thin uppercase (e.g., Futura PT Light).”
B.5 Additional Long-Form Results
Figure 27: CRAFT results for long_prompt2.
Figure 28: CRAFT results for long_prompt3.
Figure 29: CRAFT results for long_prompt4.
Appendix C Cost Comparison Prompts
C.1 cost_prompt1
“Portrait Pic, mysterious silhouette of a majestic wolf standing on a cliff at sunset, created with a dreamy double exposure effect showing a misty forest inside its silhouette, combined with expressive oil brushstrokes for texture and depth, cinematic lighting, surreal atmosphere, amazing depth, double exposure, surreal, intricately detailed, perfect balanced, deep fine borders, artistic photorealism, highly detailed, masterpiece, award winning, double exposure.”
C.2 cost_prompt2
“Generate a series of six candid, documentary-style photos of this Indonesian president in office, in the rice fields, and partying with other presidents.”
C.3 cost_prompt3
“Leela replaces the original heroine: tall, athletic cyclops woman, purple ponytail, white tank top, black trousers, shoulder harness, sweat-stained and battle-worn. She’s carrying Nibbler in her left arm: small black three-eyed alien in red cape and white shorts, clinging to her with worried expression. Leela grips a detailed sci-fi pulse rifle in one hand, matching the exact shape and silhouette of the Aliens weapon, but drawn in Futurama’s clean cartoon linework and limited shading. Blue backlight haze, drifting smoke, cinematic composition, dramatic rim-lighting on Leela’s face. Keep Leela and Nibbler true to Futurama’s animation style, smooth outlines, bold colours, simple cel-shading. Background stays faithful to the original Aliens reference: industrial sci-fi corridor, diffused fog, blue-grey tones, subtle reflections. Use both reference images for design accuracy: the pose, framing, weapon and atmosphere from the Aliens still, and Leela + Nibbler’s exact character designs from Futurama.”
C.4 cost_prompt4
“Create a single cinematic illustration that visually represents the following poem, capturing its emotions, metaphors, and atmosphere Good night mother, Good night father, Kiss your little son. Good night sister, Good night brother, Good night everyone.”
C.5 cost_prompt5
“A hyperrealistic photo of glass-made samgyeopsal placed on a grill.”
C.6 cost_prompt6
“Ultra-realistic Roronoa Zoro portrait.”
Appendix D Additional Qualitative Results
This section provides additional qualitative examples for models evaluated in Section 8.7, demonstrating further improvements achieved by CRAFT across diverse prompts and backbones.
D.1 Qwen Image Additional Examples
Figure 30: Baseline vs CRAFT. Prompt: a subway train in an empty station.
Figure 31: Baseline vs CRAFT. Prompt: a plate that has no bananas on it. There is a glass without orange juice next to it.
Figure 32: Baseline vs CRAFT. Prompt: the eye of the planet Jupiter.
D.2 FLUX-Dev Additional Examples
Figure 33: Baseline vs CRAFT. Prompt: sky above person. sky above street. sky above tree. sky above traffic light. sky above box. box on street.
Figure 34: Baseline vs CRAFT. Prompt: studio shot of vines in the shape of text knowledge is power sprouting, centered.
D.3 Z-Image Additional Examples
Figure 35: Baseline vs CRAFT. Prompt: Ultra-realistic Roronoa Zoro portrait.
Figure 36: Baseline vs CRAFT. Prompt: funny anatomical diagram of dog pet, with humorous annotations.
D.4 FLUX-2 Pro Additional Examples
Figure 37: Baseline vs CRAFT. Prompt: Infographic about the Jackson Laboratory en.
Figure 38: Baseline vs CRAFT. Prompt: Write as many cute handwritten notes and highlights as possible on the photo!
Appendix E Additional Cost Comparison Examples
This section provides additional cost–quality comparison examples between Z-Image-Turbo baseline, Z-Image-Turbo with CRAFT, and Hunyuan Image 3.0. Full prompts are available in Section C.
Figure 39: Baseline vs CRAFT vs Hunyuan Image 3 (see Appendix C, cost_prompt3).
Figure 40: Baseline vs CRAFT vs Hunyuan Image 3 (see Appendix C, cost_prompt4).
Figure 41: Baseline vs CRAFT vs Hunyuan Image 3 (see Appendix C, cost_prompt5).
Figure 42: Baseline vs CRAFT vs Hunyuan Image 3 (see Appendix C, cost_prompt6).